OpenStudy (anonymous):

simplify this expression!! 8x^4/x^3

OpenStudy (e.mccormick):

OK. So you have a fraction with an x that has an exponent on both sides. What does that mean about the exponent on the bottom?

OpenStudy (anonymous):

can you divide it into it?

OpenStudy (abb0t):

\[\frac{ a^n }{ a^m }= a^{n-m}\]

OpenStudy (e.mccormick):

Sort of, what aab0t shoed is a great example of what needs to happen here.

OpenStudy (e.mccormick):

showed... whee. I need coffee still. LOL

OpenStudy (anonymous):

the exponents are not showing on what he posted.

OpenStudy (abb0t):

\[\large \frac{ a^x }{ a^y } = a^{x-y}\]

OpenStudy (anonymous):

so it would be 8x?

OpenStudy (e.mccormick):

Yes!

OpenStudy (anonymous):

because you keep the 8.. and 4-3 is one/

OpenStudy (e.mccormick):

Exactly. What you are working on is rules of exponents. There are just a few of them and you used two at the same time. One was that an exponent on the bottom of a fraction becomes negative on the top. The other is that like bases can have exponents added. Well, if you add a negative number, it means subtraction.

OpenStudy (anonymous):

can you help me with another one with exponents? if i can figure this out im pretty sure i can handle the rest of my assignments lol.

OpenStudy (e.mccormick):

Sure.

OpenStudy (abb0t):

Yes.

OpenStudy (anonymous):

(4y-3)^3 . My textbook is telling me to multiply the first two sets of the numbers, but is not really clear on what to do with the third set of the 4y-3.

OpenStudy (e.mccormick):

OK. You end up foiling like a square for the first two, then you multiply every term again. So it is just doing it again, but with three terms rather than two.

OpenStudy (e.mccormick):

Basically uou have: (4y-3)(4y-3)(4y-3) So do the first multiplication and what do you get?

OpenStudy (anonymous):



OpenStudy (e.mccormick):

AKA: \[\large (16y^2-24y+9)(4y-3)\]

OpenStudy (e.mccormick):

All I did that was put together the two -12y terms. Now, because of the power multiplication, I also have to keep things in ( ).

OpenStudy (anonymous):

so now do i have to FOIL those numbers now

OpenStudy (e.mccormick):

Basically. I made a refeence for you.

OpenStudy (e.mccormick):

After you multiply every term like that, you combined like terms. So any x with a power gets put together with other xes of the same power.

OpenStudy (anonymous):

and you add the exponents together?

OpenStudy (e.mccormick):

No, they keep the same exponents because it is additoon, not multiplication. What you change is the coefficents. Like when I did -12y-12y=-24y

OpenStudy (anonymous):

64y^3 -144y^2+108y-27 ?

OpenStudy (e.mccormick):

Yep! That is it!

OpenStudy (anonymous):

yay thank you! i appreciate your help. i am new to this site and usually don't come to the internet to seek answers, but my math is terrible and i'm in online courses so there really isn't help available other than the textbook!

OpenStudy (e.mccormick):

Adding variables with exponents: \(\large a^n+a^m=a^n+a^m\) because the exponents are different. \(\large a^n+a^n=2a^n\) because the exponents are the same Multiplying variables with exponents: \(\large a^n\cdot a^m=a^{n+m}\) \(\large a^n\cdot a^n=a^{2n}\)

OpenStudy (e.mccormick):

Yah, having someone explain the points you do not get is always worth it!

OpenStudy (anonymous):

have you been on this site for a long time? is this all there is to do, just type your questions in?

OpenStudy (e.mccormick):

I have only been here a short while, but that seems to be the makjority of it. Learn and help other learn. However, tutoring another person is one of the best ways to force yourself to remember things. So if you help the people you can help, it makes you stronger at math. Other sites are better for lectures, like Khan Academy for individual topics. And on the Youtube Education section there are entire classes you can watch. Great to be able to pause and rewind a lecture so you do not miss anything!

OpenStudy (anonymous):